# Observing network dynamics through sentinel nodes

### v1

@@ Acknowledgements
The LFR benchmark model produces networks that have both heterogeneous degree distributions and community structure with heterogeneous community sizes [72]. We set the expected power-law exponents to $-3$ and $-1.5$ for the degree distribution and the distribution of community sizes, respectively. We set the probability of connecting nodes between communities to 0.1, the expected average degree to 4, and the minimum community size to 20. The final network had $N=998$, $M=1,988$, $\overline{k}=3.98$, and CV of the degree equal to $0.66$.
Appendix S2 Approximation error as a function of $n$
In the main text (section II.2), we showed that the degree distributions of optimized node sets remain distinct from those of completely random node sets up to $n=12$ for the dolphin and BA networks. Figure S6 indicates that the approximation errors of optimized node sets also remain distinct from and lower than those of completely random node sets. We note that the approximation error of both optimized and completely random node sets decreases as $n$ increases, which is expected.
Appendix S3 Distribution of approximation errors for various networks
@@
In Figs. 2(a) and (b) in the main text, we demonstrated that the state averaged over the nodes in the optimized node set can closely approximate the state averaged over all the nodes in the coupled double-well dynamics on the dolphin and BA networks, respectively. We show the corresponding results for all ten networks in Fig. S7, as well as for the mutualistic interaction, SIS, and gene-regulatory dynamics in Figs. S8, S9, and S10, respectively. Figures S7(a) and (g) are identical to Figs. 2(a) and (b), respectively.
@@
Across all dynamics and networks, the optimized node sets have relatively small approximation error. However, the relationship between degree-preserving and completely random node sets differs. In some networks, the difference between the degree-preserving and completely random node sets is large, such as for the dolphin network (Fig. S7a). However, it is not the case for the road network (Fig. S7d) and the LFR network (Fig. S7j). This latter result may be because the degree sequence may not contain sufficient information to characterize these networks: the road network has a complicated community structure imposed by geography [64], and the LFR network has a marked community structure by design [72]. Either of these cases may make particular nodes more suitable or less so for inclusion in $S$ for reasons other than the node’s degree. In addition, the road network has a narrow degree distribution, probably making the degree a less important indicator of a node’s importance in representing the dynamics of the entire network. Even in these cases, our algorithm finds node sets with relatively low approximation error.
Figure S7: Approximation error for the coupled double-well dynamics for node sets with $n=\lfloor\ln N\rfloor$ nodes. We compare among the optimized node sets (shown by the blue circles), degree-preserving random node sets (orange triangles), and completely random node sets (green squares). (a) Dolphin. (b) Metabolic. (c) Proximity. (d) Road. (e) Email. (f) ER. (g) BA. (h) HK. (i) GKK. (j) LFR. Panels (a) and (g) are also shown in Fig. 2 in the main text.
Figure S8: Approximation error for the mutualistic interaction dynamics. See the caption of Fig. S7 for the legends.
Figure S9: Approximation error for the SIS dynamics. See the caption of Fig. S7 for the legends.
Figure S10: Approximation error for the gene-regulatory dynamics. See the caption of Fig. S7 for the legends.
Appendix S4 Statistical results for the effect of node set type on approximation error
To systematically assess the performance of our optimization algorithm across different dynamics and networks, we generated 100 optimized node sets by running the optimization algorithm 100 times, for each combination of dynamics and network. For comparison, we also generated 100 degree-preserving node sets, one per optimized node set, 100 completely random node sets for each network, and 100 node sets from each heuristic algorithm introduced in section S6 ($k$-constrained, $k$-quantiled, $k_{\rm nn}$-constrained, and community-based) for each network. Note that neither the completely random nor heuristic algorithm node sets depend on the dynamics.
@@
We computed the approximation error for each node set obtained by each of the seven algorithms, each of the four dynamics, and each of the ten networks. To compare the approximation error, we conducted a multi-way analysis of variance (ANOVA) with three independent variables, i.e., dynamics (reference: coupled double-well dynamics), network (reference: BA network), and node set type (reference: completely random node sets). We exclude the ER network because all node sets achieve small approximation errors on this network, probably due to limited heterogeneity in the degree distribution. The dependent, or response, variable in the model is the approximation error, $\varepsilon$. We use $\ln\varepsilon$ because tiny (i.e., near zero) error values are present; the error values obey skewed distributions for most dynamics, networks, and node set types; and the variance of $\varepsilon$ tends to be large when the mean of $\varepsilon$ is large.
@@
Overall, our ANOVA model predicts $\ln\varepsilon$ well ($R^{2}=0.70$). In part due to the large sample size (i.e., 25,200 observations), all three independent variables are highly significant (dynamics: $df=3$, $F=7017.89$, $p<10^{-7}$, where the $F$ statistic is the mean-square error of the factor divided by the mean-square error of the residuals; network: $df=8$, $F=175.08$, $p<10^{-7}$; node set type: $df=6$, $F=6091.11$, $p<10^{-7}$).
@@
We show the differences between the average error for each pair of node set types, as computed by a Tukey’s honestly significant difference test, in Table S1. Each row of Table S1 specifies the estimated difference in average $\ln\varepsilon$ between the two node set types, the 95% confidence interval of the estimated mean difference, and a $p$-value for the difference adjusted for multiple comparisons. For example, the first row states that, across dynamics and networks, optimized node sets have an average $\ln\varepsilon$ value that is 5.625 smaller than that of completely random node sets. On the natural scale, the average $\varepsilon$ for optimized node sets is $277.31/e^{-5.625}=$ 277.3 times smaller than the average $\varepsilon$ for completely random node sets.
@@
Table S1: Differences in average $\ln\varepsilon$ between different node set types when we use the test dynamics for optimization. The differences are computed by a Tukey’s honestly significant difference test. The 95% confidence intervals (CI) of the differences and associated $p$-values, adjusted for multiple comparisons, are also shown.

### v3

@@ Acknowledgments
The LFR benchmark model produces networks that have both heterogeneous degree distributions and community structure with heterogeneous community sizes [84]. We set the expected power-law exponents to 3-3 and 1.5-1.5 for the degree distribution and the distribution of community sizes, respectively. We set the probability of connecting nodes between communities to 0.1, the expected average degree to 4, and the minimum community size to 20. The final network had $N=998$, $M=1,988$, $\overline{k}=3.98$, and CV of the degree equal to $0.66$.
Appendix S2 Comparison between the sentinel node approximation and $\overline{x}$ for all dynamics and networks
We show in Fig. S7 the comparison between our sentinel node approximation with $n=\lfloor\ln N\rfloor$ nodes and $\overline{x}$ for each pair of the four dynamics and 20 networks. We find that the sentinel node approximation captures the main bifurcations of the original dynamical system fairly well in all cases.
Figure S7: Comparison between the sentinel node approximation and $\overline{x}$ for each pair of dynamics and network. The sentinel node approximation using $n=\lfloor\ln N\rfloor$ nodes is shown in purple. The black lines represent the average equilibrium state of the nodes, $\overline{x}$.
Figure S7: (Continued.)
Figure S7: (Continued.)
Figure S7: (Continued.)
Appendix S3 GBB, DART, and their variants
Two known methods for dynamics reduction on networks are the Gao-Barzel-Barabási (GBB) reduction [9, 26, 15] and the dynamics approximation reduction technique (DART) [10, 11, 13, 27, 14]. These two methods attempt to match a different weighted sum of $x_{1}$, $\ldots$, $x_{N}$ from $\overline{x}$. Specifically, the observables of the GBB and DART are in the form $\sum_{i=1}^{N}a_{i}x_{i}$. For undirected networks, which we assume, $a_{i}$ is proportional to the degree of the $i$th node in the case of the GBB and the $i$th entry of the leading eigenvector of the adjacency matrix in the case of the DART.
Using the dolphin network as an example (see Fig. 1 in the main text), we show the results obtained from the GBB and one-dimensional DART approximations in Fig. S8a and S8b, respectively. The GBB approximation is faster than our algorithm to determine $S$, and the same holds true for the one-dimensional DART at least for sparse networks (see section S17). However, even with just $n=1$ sentinel node, our method performs better than the GBB and DART at approximating $\overline{x}=\sum_{i=1}^{N}x_{i}/N$, reducing $\varepsilon$ by more than 75% (GBB: $\varepsilon=0.171$; DART: $\varepsilon=0.142$). The different performance is mainly owing to the fact that, while GBB and DART track the collective state of the system (i.e., $\langle x\rangle$ in the Introduction section of the main text; brown and dark yellow lines in Fig. S8) and thus predict a single transition point, our method can discern the multistage nature of the actual observed transition (black solid line).
Figure S8: Alternative approximations to $\overline{x}$. The approximations for the GBB and DART methods are shown in red and pink, respectively. The brown and dark yellow lines show the observables (i.e., a particular weighted average of $\{x_{1}^{*},\ldots,x_{N}^{*}\}$) that the GBB and DART schemes intend to approximate. The $x_{i}^{*}$ values are shown in gray. The unweighted average $\overline{x}$ is shown in black. (a) GBB. (b) One-dimensional DART.
Our method also outperforms a variant of the GBB [12] and a multi-dimensional version of DART [10]. First, we consider a variant of the GBB that uses the Chebyshev polynomials [12], which we call the Chebyshev reduction. Consider a class of dynamical systems on networks given by
$\frac{dx_{i}}{dt}=F(x_{i})+\sum_{j=1}^{N}A_{ij}G(x_{i},x_{j}).$ (S10)
The four dynamical system models considered in the present study (i.e., coupled double-well, mutualistic species, SIS, and gene-regulatory) belong to this class. The original GBB [9] for undirected networks is given by
$\frac{d\langle x\rangle}{dt}=F(\langle x\rangle)+\beta G(\langle x\rangle,\langle x\rangle),$ (S11)
where $\langle x\rangle=\sum_{i=1}^{N}a_{i}x_{i}$, $a_{i}=k_{i}x_{i}/\sum_{\ell=1}^{N}k_{\ell}$, $\forall i$, $\beta=\sum_{\ell=1}^{N}k_{\ell}^{2}/\sum_{\ell^{\prime}=1}^{N}k_{\ell^{\prime}}$. The Chebyshev reduction approximates $F(\langle x\rangle)$ and $G(\langle x\rangle,\langle x\rangle)$ in Eq. (S11) by Chebyshev polynomials of a given order. However, it should be noted that $G(\langle x\rangle,\langle x\rangle)$ is approximated through two-dimensional Chebyshev polynomial approximation to $G(x,y)$, not simply by fitting a one-variable Chebyshev polynomial of $x$ to $G(x,x)$.
When the original $F$ or $G$ is a polynomial of the order less than or equal to that of the Chebyshev polynomials used, the Chebyshev reduction is exact. Therefore, the Chebyshev reduction of order up to three and that of order up to two are equivalent to the original GBB for the coupled double-well and SIS dynamics, respectively, for which both $F$ and $G$ are polynomials. Given this, we only demonstrate the Chebyshev reduction for the mutualistic interaction and gene-regulatory dynamics. These dynamics have a non-polynomial $G$, which we approximate by the Chebyshev polynomials. In contrast, $F$ is polynomial in these dynamics and does not need to be approximated. We use the HK network for the mutualistic interaction dynamics and the FlyBi network for the gene-regulatory dynamics. To fit Chebyshev polynomials to $G$, we need to specify the range of $x$ that we linearly transform into [1,1][-1,1]. We set this range to $0,x_{\max}]$, where $x_{\max}=26.87$ and $x_{\max}=124.58$ for the HK and FlyBi networks, respectively. With these $x_{\max}$ values, the interval $0,x_{\max}]$ represents and covers the ranges of $x_{i}$ that we observe in numerical simulations as we vary the control parameter.
We show the numerical results for the mutualistic interaction dynamics on the HK network in Fig. S9. Figure S9a–d shows that our sentinel node approximation works progressively better with a larger number of the sentinel nodes, $n$, as we showed in Fig. 1c–f in the main text for the coupled double-well dynamics on the dolphin network. Figure S9e shows the behavior of the GBB and Chebyshev reductions. We used the Chebyshev polynomials of order up to 2 and 3 for each variable (i.e., each of $x$ and $y$ in $G(x,y)$). We find that the Chebyshev reduction does not improve over the GBB reduction in approximating either $\overline{x}$ or the target of the GBB reduction, $\langle x\rangle$. The results are similar for the gene-regulatory dynamics on the FlyBi network, as shown in Fig. S10e. These results are intuitive because, in the present modeling context, the Chebyshev reduction is an approximation to the GBB reduction, whose goal is not to approximate $\overline{x}$.
Figure S9: Approximations of the average activity, $\overline{x}$, of the mutualistic interaction dynamics on the HK network. The gray lines indicate $x_{i}^{*}$ at each value of the control parameter. The black lines indicate $\overline{x}$. Note that the gray and black lines are the same in all the panels. The purple lines in (a)–(d) show the sentinel node approximation with the selected sentinel nodes, whose $x_{i}^{*}$ values are shown in orange (a), yellow (b), green (c), and blue (d). (e) GBB and Chebyshev reductions. (f) DART of different dimensions. The brown line in (e) and the dark yellow line in (f) show the observables (i.e., a particular weighted average of $\{x_{1}^{*},\ldots,x_{N}^{*}\}$) that the GBB and DART schemes, respectively, intend to approximate. In (d), the purple line almost completely covers the black line.
Figure S10: Approximations of the average activity, $\overline{x}$, of the gene-regulatory dynamics on the FlyBi network. See the caption of Fig. S9 for the legends. In (a), the orange line covers the black and purple lines. In (a)–(d), the purple line covers the black line. In (f), the thick pink line almost completely covers the dark yellow line.
We now assess a multidimensional version of DART given by Eqs. (30)–(32) in [10]. The one-dimensional DART aims to approximate the dynamics of $\langle x\rangle$ and uses the leading eigenvector of the transposed adjacency matrix, $A^{\top}$, as $(a_{1},\ldots,a_{N})$ for a one-dimensional dynamics similar to Eq. (S11[10, 11]. The $d$-dimensional DART consists in separately building the one-dimensional DART for the $d$ eigenvectors of $A^{\top}$ that are associated with the $d$ largest eigenvalues in modulus and averaging the one-dimensional approximation to $\langle x\rangle$ over the $d$ one-dimensional equations. Among the multidimensional variants of DART, we did not use the “cycle reduction” [10] because it assumes the presence of multiple eigenvalues with the same largest modulus for the adjacency matrix. An adjacency matrix would have such eigenvalues only if the network is highly symmetric or multipartite. In contrast, we are interested in general unipartite networks in the present study. We also avoided the multidimensional DART in [11] because it implicitly assumes community structure of the network, which informs the reduced adjacency or Laplacian matrix (i.e., Fig. 3 in [11], and the derivation of the $3d$-dimensional reduced dynamics for the Lorenz model on p.22 with Fig. 15). Our focus is on networks that may or may not have community structure. For the same reason, we did not use the multidimensional reduction method for networks with community structure proposed in [14].
We set $d=2$ and $d=3$ in our demonstration. We show numerical results for the mutualistic interaction dynamics on the HK network and the gene-regulatory dynamics on the FlyBi network in Fig. S9f and S10f, respectively. We find that the $d$-dimensional DART with a larger $d$ is more accurate at approximating $\overline{x}$. However, our sentinel node approximation provides a much more accurate approximation even with $n=1$ than the DART with $d\in\{1,2,3\}$, endorsing the fact that goal of DART is to approximate a specific $\langle x\rangle$, not $\overline{x}$.
To be quantitative, we calculated the approximation error in these numerical results. The approximation error against $\overline{x}$ in the case of mutualistic interaction dynamics on the HK network, shown in Fig. S9, is equal to $\varepsilon=1.35\times 10^{-2}$, $2.48\times 10^{-3}$, $9.16\times 10^{-4}$, and $5.20\times 10^{-4}$ for the sentinel node approximation with $n=1$, 22, 33, and $4$, respectively; $\varepsilon=2.03$ for GBB; $\varepsilon=2.23$ and $3.28$ for the Chebyshev reduction of order up to 2 and 3, respectively; $\varepsilon=3.72$, $2.06$, and $1.46$ for the DART with $d=1$, 22, and $3$, respectively. The approximation error against the preferred observables for GBB and DART, i.e., $\langle x\rangle$, in the same simulations is as follows: $\varepsilon=0.897$ for GBB; $\varepsilon=0.725$ and $2.43$ for the Chebyshev reduction of order up to 2 and 3, respectively; $\varepsilon=1.42$, $0.657$, and $0.535$ for the DART with $d=1$, 22, and $3$, respectively. In the case of the gene-regulatory dynamics on the FlyBi network, shown in Fig. S10, the approximation error against $\overline{x}$ is: $\varepsilon=1.28\times 10^{-3}$, $3.87\times 10^{-5}$, $1.42\times 10^{-5}$, and $2.50\times 10^{-5}$ for the sentinel node approximation with $n=1$, 22, 33, and $4$, respectively; $\varepsilon=48.9$ for GBB; $\varepsilon=666$ and $2,530$ for the Chebyshev reduction of order up to 2 and 3, respectively; $\varepsilon=55.0$, $30.4$, and $7.38$ for the DART with $d=1$, 22, and $3$, respectively. For the same simulations, the approximation error against $\langle x\rangle$ is equal to: $\varepsilon=0.213$ for GBB; $\varepsilon=102$ and $489$ for the Chebyshev reduction of order up to 2 and 3, respectively; $\varepsilon=0.0741$, $0.0562$, and $0.985$ for the DART with $d=1$, 22, and $3$, respectively. These $\varepsilon$ values are consistent with our main conclusion that the Chebyshev reduction of order up to 2 and 3 or the DART with $d=2$ and $d=3$ does not provide an accurate approximation to $\overline{x}$ compared to the sentinel node approximation.
Appendix S4 Distribution of approximation errors for various dynamics and networks
@@
In Figs. 2a and b in the main text, we demonstrated that the state averaged over the nodes in the optimized node set can closely approximate the state averaged over all the nodes in the coupled double-well dynamics and the SIS dynamics on all 20 undirected and unweighted networks. We show the corresponding results for the mutualistic species and gene-regulatory dynamics in Fig. S11a and b, respectively.
@@
Across the two dynamics and all networks, the optimized node sets have relatively small approximation error. This result is similar to that for the coupled double-well and SIS dynamics shown in Fig. 2 in the main text. However, the relationship between degree-preserving and completely random node sets differs. In particular, while the difference between the degree-preserving and completely random node sets is large for many networks, it is small for the road network and the LFR network. This result may be because the degree sequence may not contain sufficient information to characterize these networks: the road network has a complicated community structure imposed by geography [65], and the LFR network has a marked community structure by design [84]. Either of these cases may make particular nodes more suitable or less so for inclusion in $S$ for reasons other than the node’s degree. In addition, the road network has a narrow degree distribution, probably making the degree a less important indicator of a node’s importance in representing the dynamics of the entire network. Even in these cases, our algorithm finds node sets with relatively low approximation error.
@@
To formally assess the differences between node set types in terms of approximation error, we generated 100 optimized, degree-preserving, and completely random node sets for each combination of dynamics and network. We computed the approximation error for each node set. We excluded the ER network from this analysis due to the tendency to obtain tiny approximation errors regardless of the node set type. We then conducted a multi-way analysis of variance (ANOVA) with three independent variables: dynamics (reference: coupled double-well dynamics), network (reference: BA network), and node set type (reference: completely random node sets). The dependent variable is the approximation error, $\varepsilon$. We use $\ln\varepsilon$ because tiny (i.e., near zero) error values are present; the error values obey skewed distributions for most dynamics, networks, and node set types; and the variance of $\varepsilon$ tends to be large when the mean of $\varepsilon$ is large.
@@
Our ANOVA model predicts $\ln\varepsilon$ well ($R^{2}=0.76$). All independent variables are statistically significant (dynamics: $df=3$, $F=6,827.3$, $p<10^{-7}$; network: $df=18$, $F=179.81$, $p<10^{-7}$; node set type: $df=2$, $F=24,770$, $p<10^{-7}$), partly due to a large sample size.
@@
We show the differences between the average error for each pair of node set types, as computed by a Tukey’s honestly significant difference test, in Table S1. Each row of the table specifies the estimated difference in average $\ln\varepsilon$ between the two node set types, the 95% confidence interval of the estimated mean difference, and a $p$ value for the difference adjusted for multiple comparisons. For example, the first row states that, across dynamics and networks, optimized node sets have an average $\ln\varepsilon$ value that is 6.892 smaller than that of completely random node sets. On the natural scale, the average $\varepsilon$ for optimized node sets is $1/e^{-6.892}=984.4$ times smaller than the average $\varepsilon$ for completely random node sets.
Figure S11: Approximation error, $\varepsilon$, for two other dynamics on 20 undirected unweighted networks. (a) Mutualistic species dynamics. (b) Gene-regulatory dynamics. For each pair of dynamics and network, we compare three node set types (optimized: blue circles, degree-preserving: orange triangles, completely random: green squares).
@@
Table S1: Differences in average $\ln\varepsilon$ between three node set types across dynamics and networks. The differences are computed by a Tukey’s honestly significant difference test. The 95% confidence intervals (CI) of the differences and associated $p$ values, adjusted for multiple comparisons, are also shown.